TERAHERTZ AND FIR SYNCROTRON SPECTRA OF METHYL MERCAPTAN, CH_{3}SH

\underline{LI-HONG XU}, R. M. LEES, Centre for Laser, Atomic and Molecular Sciences (CLAMS), Department of Physics, University of New Brunswick, Saint John, N.B., Canada E2L 4L5; H. S. P. MÜLLER, C. P. ENDRES, O. BAUM, F. LEWEN, S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln; K. M. MENTEN, MPIfR, 53121 Bonn, Germany.

Methyl mercaptan, CH_{3}SH, is an abundant molecule e.g. in the high-mass star-forming region Sagittarius B2^{a}^{b}. It is the sulfur analog of methanol and features also threefold internal rotation of the methyl group. The most recent analysis^{c} consisted mostly of a-type transitions, treated A and E torsional species separately, and does not yield predictions of sufficient accuracy for the needs of the recently launched Herschel mission or the upcoming SOFIA and ALMA. New terahertz spectra have been recorded with microwave accuracy at the Universität zu Köln which cover thus far large portions of the 1.1−1.5 THz region as well as a section near 1.9 THz. Numerous v_{t} = 0, 1 and 2 pure rotational transitions have been assigned. FIR synchrotron spectra have recently been measured at the Canadian Light Source (CLS) between 55 and 550 cm^{−1} at 0.001 cm^{−1} resolution. 53, 15, and 2 subbands have been assigned thus far for the v_{t} = 1 − 0, 2 − 1 and 2 − 0 torsional bands along with further pure rotational transitions. We will present up-to-date results from our analysis which has been started.

aLinke et al., Astrophys. J. 234 (1979) L139

bNummelin et al., Astrophys. J. Suppl. Ser. 117 (1998) 427

cBettens et al., Astrophys. J. 510 (1999) 789